A delivery van is travelling at a constant speed of 15ms^-1 when the driver sees a pedestrian crossing 50m ahead of him. He takes exactly 1 sec before he apploes the brakes as hard as he can. It takes 3 sec for him to stop.


Will the van be able to stop before he reaches the pedestrian crossing?


1
Expert's answer
2022-03-03T14:23:46-0500

Using the kinematic formula, find the distance traveled by the van. In the first second, the motion was with constant speed, hence the distance is:


"d_1 = 15m/s\cdot 1s = 15m"

During the next 2s the motion was with constant acceleration. Thus, the distance is:


"d_2 = \dfrac{v+v_0}{2}t = \dfrac{15m/s\cdot 2s}{2} = 15m"

where "v =0 m\/s" is the final speed, "v_0 = 15m\/s" is the initial speed, and "t = 2s".

Thus, the total distance:


"d = 30m"

and the van will be able to stop before he reaches the pedestrian crossing.